# A General Infrastructure and Workflow for Quadrotor Deep Reinforcement Learning and Reality Deployment

### v1

@@ I Introduction
Our main contributions are not only the proposed platform but also a physical quadrotor DRL benchmark, discussion for training and reward mechanism, and key factors that affect sim-to-real performance:
  • We propose a comprehensive and thorough open-source solution involving hardware, firmware, and software toolkits that serve sim-to-real of quadrotor learning. Different from other works, our system is built on more commonly used PX4 autopilot, providing toolkits that allow for secondary distribution.
  • We provide a benchmark for physical quadrotor DRL on AirGym, including different types of tasks: hovering, balloon, avoid, tracking, and planning, which demonstrate the effectiveness of learning-based methods on the proposed system, even in the wild without any auxiliary positioning system.
  • We summarize key insights for DRL training tasks designs that were not easy to implement before-using Ego-centric observation to replace the global states, and other settings like waypoints sequence play in tracking task and euclidean signed distance function in planning task. Besides, key factors of transferring a policy to reality outside the laboratory are also discussed.
It is worth mentioning that the proposed toolkits are independent of each other and can be used separately as needed. Our work stands as a significant infrastructure for aerial robot learning which is crucial but was rarely comprehensively mentioned in detail in previous works, providing an open-sourced option for the community.
II Related Works
We list the previous related brilliant works in the drone robot learning field in Table.I. Herein, we conduct a comprehensive comparison and present the key insight of our work: integrating the general technical details of drone robot learning into a unified framework to establish a complete workflow, thereby accelerating research progress in this field.
II-A Sim-to-Real of Learning-based Quadrotors
Sim-to-Real is a concept and methodology commonly used in robotics, which involves developing, training, and testing algorithms or models in a simulated environment to eventually transfer and apply these algorithms or models to the real world effectively and safely [9]. The differences between simulation and reality, also known as Embodied Mismatch (EM)[19], are the main challenges for transferring a trained robot to practice. This mismatch mainly derives from several aspects[19]: (i) the deviation of the modeling during the training and testing phases; (ii) tiny differences between robots in training and testing.
For end-to-end learning policies applied on quadrotors, this mismatch is manifested in the discrepancies between the dynamics models, controller biases and visual inputs, especially in tasks that require agile 3D motion and high-speed flight. Imitation learning is an end-to-end algorithm that learns by mimicking expert policies, facilitating Sim-to-Real transfer through estimating depth images from stereo inputs, which minimize domain shift [23]. Visual scene randomization is also used to create diverse simulations, improving the robustness and real-world deployment [22]. The end-to-end deep reinforcement learning requires extensive simulation training before transferring to the real world. [40] addresses this by improving simulation data fidelity, while [12] achieves Sim-to-Real through fine-tuning with real-world data. Additionally, zero-shot transfer of RL policies, aided by accurate positioning and physical parameter randomization [39], can narrow the reality gap, and RL policies can be deployed across multiple drones using only vector state input [4].
II-B Simulator and Systems for Quadrotor Sim-to-Real
As one of the earliest works in visual drone simulators, RotorS[33] is built on the Gazebo and provides ROS interfaces for external control. Furthermore, some simulators are built on Unreal Engine and Unity like AirSim[35], FlightGoggles[7], and Flightmare[38], that offer highly realistic rendering capabilities. Flightmare is a famous flight simulator that can flexibly decouple the physics engine and renderer to greatly improve the RL simulation speed. Moreover, Flightmare uses efficient C++ implementation but at the cost of being inflexible to extend[47]. Pybullet-Drones[30] aim to serve for RL of a swarm. CrazyS[36] and CrazySim[20] are built for Crazyflie and have no RL interfaces.
TABLE I: Comparison of previous works in this aera.
Work Release Type environment Device Engine VecEnv prop rate atti vel pos Controller
Flightmare[38] 2020 simulator Flexible CPU Virtual PID
gym-pybullet-drones[30] 2021 environment custom pyBullet CPU Virtual PID
Learning high-speed flight in the wild[23] 2021 algorithm Agilicious MPC
Agilicious[5] 2022 hard/software custom Betaflight
OmniDrones[47] 2023 environment custom PhysX GPU Virtual PID
@@ II-B Simulator and Systems for Quadrotor Sim-to-Real
Champion-level drone racing using deep reinforcement learning[12] 2023 algorithm custom Agilicious CPU Betaflight
Aerial Gym Simulator[15] 2023 environment custom PhysX GPU Virtual PID
Back to Newton’s law[48] 2024 algorithm custom custom CPU Betaflight/PX4
What matters RL sim2real[3] 2024 algorithm OmniDrones Crazyfile PhysX GPU Crazyfile Controller
AirGym(Ours) 2024 hard/software/environment custom X152b PhysX CPU/GPU PX4/rlPx4controller
Thanks to the rapid development of accelerated computing, recent years have seen the emergence of many simulators capable of utilizing GPUs for parallel RL sampling to speed up computations[27]. Upgrades in simulators have significantly accelerated the development of embodied intelligent robots. The latest two contributions to GPU-vectorized drones simulator are Aerial Gym[15] and OmniDrones[47]. Aerial Gym is built on the IsaacGym and serves as the foundation for this work, upon which we make numerous significant upgrades. OmniDrones provides a platform based on Omniverse IsaacSim, particularly for drone RL tasks, containing hover, track, flythrough, and transportation tasks, etc. However, the use of Aerial Gym and OmniDrones is not aimed at transferring to real-world scenarios, but rather, they serve as a benchmark for validating algorithms. Agilicious[5] is an open-sourced system for quadrotor sim-to-real that uses a rewritten BetaFlight for low-level control. However, Agilicious imposes strict restrictions on the way its software is open-sourced and its licensing agreement, making it impossible to redistribute.
III Preliminary
III-A Problem Formulation for Aerial Robot Learning

### v2

@@ 2.2 Infrastructure & Workflow
  • Quadrotor platform X152b: an open-sourced onboard sensing and inference flight platform, shown in Fig.2, where we integrate computation and perception devices and provide the mechanical model for simulation alignment.
Table 1: Time consumption for training a policy: training from scratch.
Task Name hovering tracking avoidance target hitting planning
Comsumption (min) 2.4465 $\pm$ 0.1878 8.1619 $\pm$ 0.3811 5.3834 $\pm$ 0.4229 3.8277 $\pm$ 0.2125 7.1009 $\pm$ 0.3964
Our workflow demonstrates robust and rapid policy training-to-deployment capabilities, achieving mission-ready performance within minutes. As shown in Tab.1, the time consumptions of training a task in simulation among 10 trials are listed. In practice, the trained model is directly transferred to practical devices to realize zero-shot sim-to-real. Our training runs on a workstation with Intel Core i9-13900K processor and an NVIDIA RTX4090 graphic card, and the policy is deployed to an onboard computer with RK3588s processor. Since there is no collision during the training and flying, we can simply copy policies from the simulator to real-world robots without any fine-tuning which has been proved to be feasible in previous studies[5]. Additionally, our efficient workflow pipeline software offers one-click solutions for complex engineering implementations like robot state estimation and perception pre-processing, enabling immediate deployment and inference. Furthermore, the workflow is highly scalable; the modules and tools within the pipeline are not limited to the specific hardware platform we used. Features and usage of subsystems are described in supplementary materials.
Figure 2: The proposed pipeline and the relationships between each part. In (a) AirGym, we provide four classical tasks: tracking, avoidance, target hitting, and planning. (b) illustrates the sensing processing and features fusion during the training phase. The left image in (b) shows visual input is compressed into a vector using VAEs. The depth image is reconstructed by using self-supervised learning, and the encoding part is used as the image encoder in the DRL loop. The right image in (b) illustrates ESDF distance which is used as an efficient spatial description, where we use the minimum value in depth array is used as the ESDF distance. (c) shows all different control modes that can be selected. (d) is the onboard inference module that runs on the flight platform. It is worth mentioning that the VIO part is not indispensable in our experiments because global states estimation is not necessary in our Ego-centric design. (e) builds the bridge between PX4 flight controller and the algorithm. Finally, bottom-left device is our flight platform X152b.
2.3 Principles in Training Design
Advances in Sim-to-Real for aerial robots increasingly enable tackling more complex tasks. However, designing effective task-oriented DRL training remains challenging. Poorly defined training objectives may misguide the learner into blind alleys, reducing the problem to unsolvable reward engineering; conversely, well-crafted training acts as a scaffold, steering learning toward generalizable and robust policies. This paradigm shift necessitates a principled approach to task design. In the following, we outline the fundamental principles to guide effective training design.
2.3.1 Control & Planning
The most fundamental question in learning-based control is deceptively simple: what exactly should the agent learn to control? The answer lies in selecting an appropriate level of abstraction for the control command—ranging from low-level motor thrusts to high-level position targets. Inspired by classical hierarchical control theory, which organizes control commands hierarchically, we adopt a similar strategy and define five distinct control modes: $PY$, LVLV, CTACTA, CTBRCTBR, and $SRT$. CTBRCTBR emerges as a robust mode, striking a balance between expressiveness and tractability. It performs consistently well across a range of tasks, indicating strong generalization capabilities. Meanwhile, $LV$ mode demonstrates faster convergence in tasks with looser precision requirements, making it a practical alternative when high-frequency control is unnecessary. Notably, the choice of control mode directly influences shapes the agent’s action space and physical embodiment, shaping both the learning space and the physical embodiment of policy execution. Thus, it is crucial to ① select appropriate control mode that matches the task at hand. For instance, while $SRT$ offers the most sensitive control, it poses significant learning challenges due to the complexity of the underlying dynamics. In contrast, mid-level abstractions such as $CTBR$ or $LV$ strike a more favorable balance between controllability and learning efficiency.
Planning complex trajectories poses significant challenges, while ② spatiotemporal awareness guidance can facilitate it. Whether through incorporating desired poses as prior knowledge that embeds useful priors into the learning process[31] or providing environmental spatial descriptors in the state space as a temporary target[5, 32], both approaches can effectively enhance flexibility when handling complex trajectories. Whether these cues come from human commands or perception pipelines, their value lies in anchoring the agent’s decision-making in both space and time, allowing the drone to plan ahead—without relying on brittle rule-based controllers. Therefore, when an agent gets lost in the "reward engineering", it is important to rethink whether sufficient guidance has been provided to the agent, and whether this guidance has been properly incorporated into the state space.
2.3.2 Environmental Sensing
In most traditional and classical quadrotor control and navigation scenarios, researchers typically use (i) the takeoff point as the origin of coordinate and (ii) adopt a North-East-Down (NED) coordinate frame. Because a locally constructed coordinate system like this allows the mathematical formulation of global/local path planning problems to be clearly defined from a third personal view, which is well-suited for control and planning approaches based on optimization but often hinders training in more embodied scenarios. The ③ egocentric sensing, where the drone perceives the world from its own point of view, enhancing learning stability and efficiency. We define an ego-centric coordinate system $\mathcal{E}$: the origin of this coordinate is fixed on the mass point of quadrotor, using the forward-facing direction of vehicle as the positive $\bm{x}_{\mathcal{E}}$. $\bm{z}_{\mathcal{E}}$ points vertically upward (toward the sky), and the $\bm{y}_{\mathcal{E}}$ is determined according to the right-hand rules. Ego-centric frame $(\bm{x}_{\mathcal{E}},\bm{y}_{\mathcal{E}},\bm{z}_{\mathcal{E}})$ is similar to the body frame $(\bm{x}_{\mathcal{B}},\bm{y}_{\mathcal{B}},\bm{z}_{\mathcal{B}})$ while the difference is $\bm{z}_{\mathcal{E}}$ in ego-centric frame always points to the sky. By anchoring perception to the drone’s forward-facing frame, the system avoids the confusion introduced by arbitrary global positions or orientations. In practice, designing the observation space with respect to the agent’s own perspective can greatly enhance learning stability and efficiency. This aligns not only with biological inspiration—humans and animals rely on self-centered perception—but also with practical benefits: observations become more compact, generalizable, and better suited for transfer to real-world deployments. Additionally, emphasize the importance of using unique state representations to ensure a one-to-one mapping from observations to actions. For example, while quaternions are commonly used for attitude representation. Since quaternions are not unique representations of the same rotation, we recommend using rotation matrices to represent orientation. This helps the neural network learn an exclusive mapping from input to output, and it makes the training converge faster.
Moreover, what a drone perceives is only as valuable as its ability to interpret it, which highlights the importance of ④ efficient perception encoding and spatial description. While raw depth images or point clouds offer rich environmental information, they are often too high-dimensional and noisy for direct use in learning policies. Instead, feature extraction—either through hand-crafted heuristics or learned encoders—can distill relevant information into compact latent variables. Probabilistic models like variational autoencoders (VAEs) are especially promising [25]. They can produce low-dimensional features that still preserve critical environmental cues. These encodings also offer smooth variations and a structured latent space, which are beneficial for stable training. When used with additional spatial priors, such as Euclidean Signed Distance Fields (ESDFs), they enable reward functions that naturally reflect obstacle proximity—encouraging safe and agile navigation [26]. Therefore, both perception encoding and spatial description are indispensable: the former enhances the effective utilization of perceptual information, while the latter bridges perception and reward design.
2.3.3 Training
DRL training for drones must address the complexity of tasks. Many tasks involve intricate logical dependencies or require the integration of multiple skills that are hard to learn simultaneously, like dodging dynamic obstacles while hovering. A practical solution is to employ ⑤ training curriculum, such as deep hierarchical learning (DHL), to decompose the task into manageable sub-problems[33]. For example, a system can first learn a perception module that maps depth inputs to compact feature representations, and then train a separate policy to act based on those features to improve training stability and overall performance.
Random environmental policy is essential for improving the generalization capability of learned policies. For example, such as tracking arbitrary trajectories, drone racing across various tracks, or navigating in unknown scenes, the ability of a model to adapt to diverse conditions is a fundamental requirement. Randomization strategies have been shown to be effective in the previous works, like environment policy in study [34] and the unified tracking policy in SimpleFlight [11].
In more complex tasks, ⑦ differential simulation-based learning yields more efficient and intelligent results[7]. This approach involves designing the RL process—especially policy updates—as an end-to-end differentiable system, enabling the entire learning process to be efficiently optimized through gradient descent. Traditional RL often relies on sampling and the design of reward functions to obtain gradients, where the optimization objectives are often non-differentiable or suffer from high gradient estimation noise. However, differentiable reinforcement learning enables direct backpropagation of gradients, leading to lower estimation noise and more stable learning dynamics[15]. Therefore, rather than designing complex reward functions, it may be more effective to consider how to construct loss functions directly from differentiable components.
2.4 Task-oriented Flying
We implement four classic RL/DRL tasks and build a set of quadrotor experiments on top of AirGym: tracking, avoidance, target hitting, and planning, as shown in Fig.2(a). All tasks support end-to-end policy training across various control levels, allowing us to select and integrate specific levels where we wish to train and maintain classical control methods. tracking is a classical drone robot learning task: the quadrotor follows the given trajectory with a desired speed. Herein we primarily conduct tracking of a lemniscate trajectory flight. Task avoidance is a new scenario not addressed in previous RL tasks: the quadrotor starts at a fixed point, then an obstacle will be initialized at a random position with an appropriate speed to hit the quadrotor. The quadrotor needs to learn a policy to avoid the obstacle and keep hovering at the original position. target hitting is a waypoint navigation task: the quadrotor flies at a very high speed, rapidly dashing toward the target. Finally, we propose a planning task: a cluster of trunks is randomly generated at the beginning of an episode, and a quadrotor needs to fly through the forest and reach the target.
In our experiments, we evaluate the results of all different control types. The $CTBR$ control mode is recommended as its effectiveness has been studied and shown to outperform others [18, 31]. Our experimental results also support this perspective: although $CTBR$ may not offer the fastest training speed or the highest reward, its flight behavior is more stable and natural, based on real-world observations. Therefore, we uniformly use the $CTBR$ mode for sim-to-real experiments.
2.4.1 Tracking with Unknown Disturbance
Figure 3: (a)(b) show the tracking performance under different conditions. We set a lemniscate trajectory as tracking reference and illustrate the cumulative flight path under indoor no-wind and outdoor windy environments. (c) is dynamic obstacle avoidance task by end-to-end DRL Sim-to-Real using depth sensing. The quadrotor with blue halo tries to dodge a fast flying football. Our experiment is conducted in a wild GNSS-deny environment and only relies on onboard sensing for perception and localization. The speed of flying ball reaches about $15\mathrm{m\cdot s^{-1}}$. (d) shows the generalization experiments by using different throwing objects with various speed. One hundred episodes are evaluated for each experiment and success rate are recorded.
Tracking a lemniscate is a commonly used baseline for demonstrating control performance. Here, we conduct lemniscate tracking experiments under both ideal indoor and dim, windy outdoor conditions, as shown in Fig.3(a)(b) respectively. Meanwhile, we present the tracking performance at two different tracking speeds.
Table 2: Comparison between different conditions of lemniscate tracking.
Items Speed Relative Tracking Errors Success Rate
SimpleFlight (Baseline) 1.6m/s 0.028 $\pm$ 0.0
2.5m/s 0.051 $\pm$ 0.002
Indoor with no wind 1.6m/s 0.022 $\pm$ 0.013 $\searrow$ 1
2.5m/s 0.03 $\pm$ 0.006 $\searrow$ 0.9
Outdoor with no wind 1.6m/s 0.023 $\pm$ 0.015 $\searrow$ 1
2.5m/s 0.032 $\pm$ 0.024 $\searrow$ 0.8
Outdoor with wind 1.6m/s 0.073 $\pm$ 0.043 0.3
2.5m/s 0.11 $\pm$ 0.041 0.2
Outdoor with wind & DR 1.6m/s 0.043 $\pm$ 0.028 $\searrow$ 0.6 $\nearrow$
2.5m/s 0.082 $\pm$ 0.035 $\searrow$ 0.5 $\nearrow$
As illustrated in Fig. 3(b), we use the Mean Euclidean Distance (MED) to evaluate the tracking errors. The results report a maximum $0.09\pm 0.07$ meters of MED under no-wind conditions. However, tracking lemniscate under moderate breeze conditions presents significant challenges. The tracking performance is shown in Fig. 3(b). Because the wind direction and strength in the outdoor environment are both uncertain, the disturbances can be considered approximately random. Thus, the trajectory is irregular, and it deviates from the reference for most of the process. The wind disturbance directly affects the quadrotor’s attitude, decreasing the control accuracy of the neural network. On the other hand, persistent wind disturbances may disrupt the quadrotor’s control, making it difficult to follow up the desired playing trajectory and leading to a crash. When delay or ahead of trajectory playback occurs, the neural network encounters states that it has never seen during training, which can lead to unpredictable outcomes [35]. To conquer this, domain randomization can significantly improve the robustness, as shown in Tab.2, by adding the temporal margin [11] and randomization on observation. Furthermore, artificially adding Gaussian-distributed external force disturbances to simulate wind as a kind of DR method can also increase the success rate. The results show that the success rate nearly doubles after DR is applied.
To make the results more persuasive, we make a comparison with SimpleFlight[11]. Due to the different scales of flight platforms, we use the relative proportion of tracking error for a fairer comparison. Note that the maximum edge length of the trajectory in SimpleFlight is limited to (1,1)(-1,1), whereas in our experiment it reaches (3,3)(-3,3). In our experiment, the relative tracking error achieves a lower average compared to the baseline, as shown in Tab. 2, which demonstrates the effectiveness of our method in outdoor environments. Moreover, we apply DR to the timing of sending the desired tracking trajectory by introducing slight delays and random time offsets $T\sim\mathcal{N}(\mu,\sigma^{2})$, where $(\mu,\sigma)=(0.3,0.5)$. Compared with the experiments without DR under windy conditions, the incorporation of DR effectively improves the success rate of the tracking experiments and reduces the relative tracking errors. It is worth mentioning that the variance of our method is higher than the baseline because the observation in our experiments comes from VIO while an external optical positioning system in SimpleFlight. Although VIO drift and quadrotor vibrations may increase the variance in tracking performance, they do not affect the overall tracking trend, as our tracking error is also computed relative to the position data from VIO.
2.4.2 Dynamic Obstacle Avoidance
The experiments above already demonstrate the system’s Sim-to-Real stability for routine tasks. In this experiment, we aim to highlight the end-to-end inference capability when faced with rapidly changing environments. As illustrated in Fig. 3(c), the quadrotor has to avoid a rapidly flying object, a ball or a cube assembled from EVA foam mats, thrown from $6$ meters away. The speed of the flying ball is $10\sim 15$ m/s. Fig.3(c) is the processed depth image from the camera, where depths beyond $4.5$ meters are truncated. Considering the high speed of the obstacle, it places high demands on the real-time neural network inference. The success rate of dynamic obstacle avoidance reaches over $90\%$ across more than 20 trials involving either a ball or a cube.
To evaluate the generalization capability of our end-to-end model, we conduct transfer experiments by training models with two different types of dynamic obstacles—ball and cube—each at distinct throwing speeds, and subsequently testing them under varying conditions. As illustrated in Fig.3(d), the training involved dynamic obstacles moving at speeds ranging from $4\;\mathrm{m\cdot s^{-1}}$ to $8\;\mathrm{m\cdot s^{-1}}$. We perform cross-testing using models trained under different configurations to assess performance when encountering previously unseen obstacle dynamics. Additionally, we increase the maximum velocity of incoming obstacles during testing to examine the robustness of the learned policies in more extreme scenarios. The results indicate that the trained model exhibits a certain level of robustness when dealing with dynamic objects of unknown types. The model trained with balls as dynamic obstacles demonstrates better generalization capability, likely because the balls move faster than cubes during training. As a result, the trained model can more easily adapt its motion planning strategy to handle high-speed objects.
2.4.3 High-speed Flight
Figure 4: High-speed target hitting results. The blue halo is the quadrotor X152b with a strip light. (a) shows the high speed tracking with a sigmoid reference. (b) shows a target hitting task that accelerates from $0\mathrm{m\cdot s^{-1}}$ and hits a virtual target.
In order to evaluate the performance of our workflow in dealing with extreme maneuvers, we conduct a series of high-speed impact and tracking experiments based on the target hitting and tracking tasks, as illustrated in Fig.4. The blue trajectory demonstrates the high-speed flight path. Here we test the sim-to-real performance on both the target hitting and tracking tasks. In the target hitting task, a virtual balloon is placed at a fixed position, and the quadrotor must accelerate from rest to strike the target. Furthermore, in the tracking task, the quadrotor follows a sigmoid-shaped trajectory starting from rest. In both experiments, no upper speed limit is imposed, allowing the quadrotor to reach the maximum velocity it can learn in AirGym. Linear and angular velocity records are shown on the right side of Fig.4. Remarkably, the platform achieves a top speed exceeding $10\;\mathrm{m\cdot s^{-1}}$, demonstrating its strong maneuverability.
2.4.4 Navigation in Forest
We set up an outdoor wild flight scenario where the quadrotor navigates through and out of a small forest, illustrated in Fig.5. Herein, we evaluate the effectiveness of policies trained in the planning task. The quadrotor is initialized at the start point and then flies to the end with $1.5\;\mathrm{m\cdot s^{-1}}$ maximum speed. The quadrotor navigates using depth images as input. Since the branches in the wild are dense and chaotic, differing somewhat from the simulation scene, we strictly constrain the actual flight altitude to around 0.5 meters to adjust the field of view (FOV) to prevent the camera from capturing images of canopies and other elements not present in the simulation. Moreover, we clip the depth image range to $0\sim 4.5$ meters to further reduce the likelihood of capturing canopies. The quadrotor autonomously navigates a distance of over $20$ meters and successfully reaches the target location as shown in the bird’s-eye view Fig.5(b). In Fig.5(c), we use the recorded depth stream to generate the occupancy map of the area, highlighting the complexity of the forest environment.
3 Discussion
The direction of robot learning has emerged as a transformative force in aerial robotics, enabling autonomous systems to adapt, learn, and perform complex tasks in dynamic environments. This transformation goes beyond the conventional approach of merely grafting traditional pattern recognition methods onto aerial robot platforms to enhance onboard capabilities. Instead, it fundamentally redefines the control, planning, and perception architecture of quadrotor from the ground up, aiming to develop one task-driven end-to-end policy to realize all functions. This trend has sparked the emergence of increasingly challenging tasks in aerial robotics, as evidenced by the surge of related research in recent years[11, 36, 19, 8, 18, 37, 4, 38, 25, 12, 27, 8, 39, 13, 15, 16, 5, 20, 10, 40, 41, 42, 32, 14, 43, 44, 9, 45, 31, 21, 34, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 22, 25, 57, 58, 59, 60, 61, 62]. In an increasing number of complex tasks, traditional methods are being replaced by end-to-end learning-based approaches, which serves as strong evidence of the generalization and effectiveness of learning-based methods. Among these, the majority of approaches are based on RL and its related variants.
Table 3: Comparison of previous highlight works in learning-based aerial robotics field.
Work Release Type Method Tasks SRT CTBR CTA LV PY/Traj Controller Principles
Control of a quadrotor with reinforcement learning[38] 2017 Methodology RL Control
FlightGoggles[21] 2019 Simulator Drone Racing
Deep drone racing[8] 2020 Methodology DL Drone Racing
Flightmare[44] 2020 Simulator RL Control / Drone Racing
NeuroBEM[63] 2021 Dataset DL Agile Flight / Control Betaflight
gym-pybullet-drones[37] 2021 Simulator RL Control / Multi-agent Lead-follower
Learning high-speed flight in the wild[39] 2021 Methodology IL Planning MPC ②③⑥
@@ 3 Discussion
Autonomous drone racing with deep reinforcement learning[5] 2021 Methodology RL Drone Racing Betaflight ③④⑥
Agilicious[27] 2022 Software Infrastructure for Any Tasks Betaflight
Aggressive quadrotor flight using curiosity-driven reinforcement learning[64] 2022 Methodology RL Planning AscTech Autopilot ④⑤
A deep multi-agent reinforcement learning framework for autonomous aerial navigation to grasping points on loads[42] 2023 Methodology RL Multi-agent Transport Crazyflie Controller
OmniDrones[31] 2023 Simulator RL Control / Tracking / Transport / Formation
Champion-level drone racing using deep reinforcement learning[4] 2023 Methodology RL Drone Racing Betaflight ③④
Reaching the limit in autonomous racing[32] 2023 Methodology RL Drone Racing Betaflight ③④⑥
Aerial gym simulator[43] 2023 Simulator RL Control / Planning
Deep-PANTHER[60] 2023 Methodology RL Planning
Learning quadrotor control from visual features using differentiable simulation[15] 2024 Methodology RL Control Betaflight
Back to Newton’s law[7] 2024 Methodology RL (Multi-agent) Planning Betaflight/PX4 ③⑥⑦
What matters RL sim2real[11] 2024 Methodology RL Control / Tracking Crazyflie Controller ①②⑥
Dashing for the golden snitch[16] 2024 Methodology RL Multi-agent Tracking Betaflight
Reinforcement learning for collision-free flight exploiting deep collision encoding[25] 2024 Methodology RL Planning PixRacer ③④⑤
Learning-based navigation and collision avoidance through reinforcement for UAVs[19] 2024 Methodology RL Planning / Avoidance
Flying on point clouds with reinforcement learning[26] 2025 Methodology RL Planning PX4 ③④⑥
NavRL[65] 2025 Methodology RL Planning PX4 ③④⑥
AirGym (Ours) 2025 Workflow RL Control / Tracking / Target Hitting / Avoidance / Planning / Customized PX4/rlPx4controller ①②③④⑤⑥
Despite the numerous learning-based methods that have been successfully implemented, a significant challenge for many newcomers to the field lies in the complexity of building and implementing such system-level engineering solutions. Moreover, the reproduction process often depends heavily on extensive engineering experience, making it difficult to achieve rapid methodological iteration and innovation. Some previous highlight aerial robot learning works are listed in Tab.3. The last column outlines the task design principles used in previous approaches which support our task-oriented framework. Comparing with these, our framework and workflow differ significantly from existing works that primarily focus on methodological innovation or simulator. Instead, we emphasize the evaluation and integration of proven or widely accepted approaches, provide a comprehensive explanation of key principles for task design, and highlight the challenges of policy transfer along with practical solutions. In conjunction with the aforementioned design principles and techniques, the simulator, software, firmware, and hardware modules we develop collectively form a complete and cohesive workflow. Based on this, our task-oriented training leverages DRL to train policies and enable zero-shot sim-to-real transfer to real-world aerial robots. We hope this work serves as a stepping stone to inspire and enable more researchers to quickly engage in practical exploration within this field, rather than being hindered by limitations in infrastructure.
Our research has also identified key limitations, especially in the sampling efficiency and generalization of DRL. Model-free RL methods often rely on vast amounts of trial-and-error interactions, making them computationally expensive and data-inefficient. Vectorized parallel simulation with simultaneous sampling in this work is one solution; however, it does not mask the inherent inefficiency of the sampling. Recently, differentiable simulations enable gradient-based optimization by backpropagating through environment dynamics, which can drastically reduce the number of required training samples[15, 7]. Although it offers a promising direction for improving sampling efficiency compared to traditional model-free RL, its advantages are not yet fully realized in complex, real-world scenarios. The applicability of differentiable simulation is currently constrained by its reliance on fully differentiable environments—a requirement that many physical or stochastic systems do not satisfy. Bridging this gap could unlock more efficient training paradigms but remains an open research challenge.
Generalization in task-oriented DRL policies remains a critical hurdle. While many DRL agents perform well on narrowly defined tasks, they often struggle when exposed to even slight variations in deployment scenarios. The most direct and effective approach to improving generalization is scaling up model parameters and training data. The success of large-scale models—such as large language models (LLMs) and vision-language models—demonstrates that scaling laws, where performance predictably improves with increased model size, data diversity, and computational resources, lead to remarkable generalization abilities. Analogously, aerial robot policies trained on massive, multi-domain datasets—combining simulation and real-world data across varying weather and obstacle conditions—are expected to exhibit stronger out-of-distribution (OOD) generalization. Embodied learning with large-scale models in aerial robotics thus offers a promising direction toward achieving robust generalization. Nonetheless, given the safety-critical and low-latency demands of aerial robot control, careful balance between model size and inference speed remains essential.
4 Methods
4.1 Deep Reinforcement Learning